OpenStudy (anonymous):

I have to write the equation of each line With y- intercept = -6 and perpendicular to y = 2x =1

OpenStudy (xkehaulanix):

Is it y = 2x = 1 or y = 2x + 1?

OpenStudy (anonymous):

=1

OpenStudy (anonymous):

x + 2y=-12

OpenStudy (anonymous):

still dont get it

OpenStudy (xkehaulanix):

...That's a weird problem. Okay then, if the equation given is y = 2x = 1, essentially it simplifies to 2x = 1 x = 1/2 This is a vertical line at x = 1/2. To get a line perpendicular to it, you would need a horizontal line with slope = 0. Taking the equation of a line to be y = mx + b, with b being the y-intercept: y = 0x - 6 y = -6 ...Are you sure you typed the problem correctly? That seems like a really weird question to give.

OpenStudy (anonymous):

Yes I did type it in correct lol